i need the correct answer

(1 pt) Calculate S/sf(x, y, z) ds For 231 where 0 Sxys 11- f(x, y, z) = Isf(x, y, z) ds = (192/5)pi Part of the surface x
Answer #1

som ? Calculate sig frey, 2) ds=2 Given surface is x= zs, osnys 112, fox, y, z)=re because, Area = Sd for 9, 2) as (2) -0)Z=1542 312 li Area=s [[Z3J17+9z+] 41. dz Zo Area=113 j2=113 Cz! J1+12+] dz Z=o dk =z²dz (11) 36 Let k=14924 dk= 36 2² dz > 1